OpenStudy (anonymous):

Help please. holly wants to save money for an emergency. Holly invests $1,500 in an account that pays an interest rate of 6.75% . How many years will it take for the account to reach $7,300? Round to the nearest hundredth

OpenStudy (anonymous):

A= P(1+r)^t is the equation

OpenStudy (anonymous):

@chris00

OpenStudy (anonymous):

@BlooooooBlahBLIIIIK

OpenStudy (anonymous):

now is the interest rate annually?

OpenStudy (anonymous):

It doesnt state if its annually or not

OpenStudy (anonymous):

\[A=P \left( 1+r \right)^{t}\] where P=$1500, r=0.0675 A=$7,300

OpenStudy (anonymous):

we need to find 't'

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Would it be 24.23??

OpenStudy (anonymous):

\[7300=1500\left( 1+0.0675 \right)^{t}\] \[7300=1500\left( 1.0675 \right)^{t}\] Dividing both sides by 1500, \[\frac{ 73 }{ 15 }=1.0675^{t}\] Taking the logarithm, \[\ln \left( \frac{ 73 }{ 15 } \right)=tln \left( 1.0675 \right)\] Making 't' the subject, \[t=\frac{ \ln \left( \frac{ 73 }{ 15 } \right) }{ \ln \left( 1.0675 \right) }\]

OpenStudy (anonymous):

which is t=24.23!

OpenStudy (anonymous):

Yay thank you for much for explaining it for me!!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

remember it is 24.23 years. never forget units!!!